Q2. Charge A has a charge of -25 HC and is 1.5 m to the left of charge B, which has a charge of +32 pC. Charge B is 1.7 m to the left of a third charge pC, which has a charge of-16 μC. If all three charges are collinear, what is the net electrostatic force on each of the following? #10-6 (a) Charge B; (b) Charge C (Read Example 10.3 &10.4 on textbook)
Answer #1


iven Coy f A Distance betwee n B(- )-7m: 202 a) Net orce on Charge B Fonce on B due to A (-35×10-6 32メ10 -3.1955 N

Force on B due to C, Fe g%. -eres (32×1016)(-16x10-6) -1 = -1.5922 N. Negative sign โ스 rfde indicating atthachí ve tone So, once o s like FA Fe = 1.6033 N l. 6033 N towrds A nekB b) Nettonce on chong-。 Fonce on C due to A , FAF - 2Sx1o6 о< -0.35 IDN (I. St 1 7)2 Rebpullive in nature

Force on dueto Bv e 2 (1门)- Net Jorce en FA Frekc = 1 1-1忌! = 1.5922-0.35)o = 1.24は net C

Note: In the final answers instead of signs the type of resulting net force is expressed as direction of attraction.